# Toeplitz Operators Defined between Köthe Spaces

### v1

@@ Abstract.
The aim of this paper is to define Toeplitz operators between Köthe spaces, especially power series spaces. We determine the conditions for continuity and compactness of these operators. We define the concept of S-tameness of a family of continuous operators. We construct the conditions on $S$-tameness for the families consisting from Toeplitz operators.
@@ 1. Introduction
The theory of Toeplitz operators defined on a Hilbert space such as Hardy and Bergman spaces is well-studied. The significance of this theory lies in the connection between operator theory and function spaces. In recent years, Toeplitz operators, whose ”associated” matrix is Toeplitz, are defined for more general topological vector spaces. For instance, in [3], Domański and Jasiczak developed the analogous theory for the space of $\mathcal{A}(\mathbb{R})$ real analytic functions on the real line. This space is not Banach space, even not a metrizable space. In [6], Jasiczak introduced and characterized the class of Toeplitz operators on the Fréchet space of all entire functions $\mathcal{O}(\mathbb{C})$. In [6], Jasiczak defined a continuous linear operator on $\mathcal{O}(\mathbb{C})$ as a Toeplitz operator if its matrix is a Toeplitz matrix. The matrix of an operator is defined with respect to the Schauder basis $(z_{n})_{n\in{N_{0}}}$. In this case, the symbol of a Toeplitz operator comes from the space $\mathcal{O}(\mathbb{C})\oplus(\mathcal{O}(\mathbb{C}))^{\prime}_{b}$, here $(\mathcal{O}(\mathbb{C}))^{\prime}_{b}$ is the strong dual of $\mathcal{O}(\mathbb{C})$. The space of entire functions $\mathcal{O}(\mathbb{C})$ is isomorphic to a power series space of infinite type $\Lambda_{\infty}(n)$. By taking inspiration from Jasiczak paper [6], we will define Toeplitz operators on more general power series spaces of finite or infinite type. Similarly, we will show that the symbol space is $\Lambda_{r}(\alpha)\oplus(\Lambda_{r}(\alpha))^{\prime}$, for $r=1,\infty$. We will also search the compactness of these operators and mention the conditions for the tameness of the families consisting of Toeplitz operators.
@@ 2. Preliminaries
In this section, we provide some fundamental facts and definitions essential for subsequent discussions. We will use the standard terminology and notation of [9]. A complete Hausdorff locally convex space E whose topology defined by countable fundamental system of seminorms $(\|\cdot\|_{k})_{k\in\mathbb{N}}$ is called a Fréchet space. A matrix $(a_{n,k})_{k,n\in\mathbb{N}}$ of non-negative numbers is called a Köthe matrix if it is satisfies that the following conditions:
@@
is called a Köthe space. Every Köthe space is a Fréchet space given by the semi-norms in its definition. By Grothendieck-Pietsch Criteria [9][Theorem 28.15] we can say that a Köthe space $K(a_{n,k})$ is called nuclear if for every $k\in\mathbb{N}$, there exists a $l>k$ so that
@@
For a nuclear Köthe spaces, $\|x\|_{k}=\sup_{n\in\mathbb{N}}|x_{n}|a_{n,k}$, $k\in\mathbb{N}$ forms an equivalent system of seminorms to the fundamental system of seminorms $\|x\|_{k}=\sum^{\infty}_{n=1}|x_{n}|a_{n,k}$, $k\in\mathbb{N}$. Dynin-Mitiagin Theorem [9][Theorem 28.12] states that if a nuclear Fréchet space E with the sequence of seminorms $(\|e_{n}\|)_{n\in\mathbb{N}}$ has a Schauder basis $(e_{n})_{n\in\mathbb{N}}$, then it is canonically isomorphic to a nuclear Köthe space defined by the matrix $(\|e_{n}\|_{k})_{n,k\in\mathbb{N}}$. So, nuclear Köthe spaces hold a significant place in theory of nuclear Fréchet space. Let $\alpha=\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ be a non-negative increasing sequence with $\lim_{n\rightarrow\infty}\alpha_{n}=\infty$. A power series space of finite type is defined by
@@
$\mathcal{O}(\mathbb{C}^{d})\cong\Lambda_{\infty}(n^{\frac{1}{d}})\quad\quad\text{and}\quad\quad\mathcal{O}(\mathbb{D}^{d})\cong\Lambda_{1}(n^{\frac{1}{d}})$
where $\mathbb{D}$ is the unit disk in $\mathbb{C}$ and $d\in\mathbb{N}$. Let $E$ and $F$ be Fréchet spaces. A linear map $T:E\to F$ is called continuous if for every $k\in\mathbb{N}$ there exists $p\in\mathbb{N}$ and $C_{k}>0$ such that
@@
for all $x\in E$. A linear map $T:E\to F$ is called compact if $T(U)$ is precompact in $F$ where $U$ is a neighborhood of zero of E. In this paper, we fixed the symbol $e_{n}$ to denote the sequence
@@
[…]
@@
    If $K(b_{n,k})$ is nuclear, then $T:K(a_{n,k})\to K(b_{n,k})$ is a compact operator if and only if there exists m such that for all k
@@
for each $x\in E$ and which define the topology. Every Fréchet space admits a grading. A graded Fréchet space is Fréchet space with a choice of grading. For more information, see [5]. In this paper, unless stated otherwise, we will study with the graded Fréchet spaces. A pair of graded Fréchet spaces $(E,F)$ is called tame provided that there exists an increasing function $\sigma:\mathbb{N}\to\mathbb{N}$, such that for every continuous linear operator T from $E$ to $F$, there exists an $N$ and $C>0$ such that
$\quad\quad\quad\quad\quad\quad\quad\quad\|Tx\|_{n}\leq C\|x\|_{\sigma(n)}\quad\quad\quad\quad\quad\quad\forall x\in E\;\;\text{and}\;\;n\geq N.$
A Fréchet space $E$ is called tame if $(E,E)$ is tame. Tameness gives us a kind of control on the continuity of the operators. In [4], Dubinsky and Vogt used this notion to find a basis in complemented subspaces of some infinite type power series spaces. We refer to [1] and [7] for other studies about tameness. In this paper, instead of all operators defined on a Fréchet space, we will restrict ourselves to a subfamily of operators. With this aim, we will give the following definition.
@@
$\quad\quad\quad\quad\quad\quad\quad\|Tx\|_{k}\leq C\|x\|_{S(k)}\quad\quad\quad\quad\text{for all}\;\;x\in E\;\;\text{and}\;\;k\geq k_{0}.$
@@
Let $K(a_{n,k})$ and $K(b_{n,k})$ be nuclear Köthe spaces and $S:\mathbb{N}\to\mathbb{N}$ be a non-decreasing function. A family of linear continuous operators $\mathcal{A}\subseteq L(K(a_{n,k}),K(b_{n,k}))$ is S - tame if and only if for every operator $T\in\mathcal{A}$ there exist $k_{0}\in\mathbb{N}$ and $C>0$ such that
$\hskip 86.72377pt\|Te_{n}\|_{k}\leq C\|e_{n}\|_{S(k)}\hskip 86.72377pt\forall n\in\mathbb{N},k\geq k_{0}.$
@@
The sufficient part is clear. To the necessary part assume that for every $T\in\mathcal{A}$ there exist a $k_{0}\in\mathbb{N}$ and $C>0$ such that
$\hskip 86.72377pt\|Te_{n}\|_{k}\leq C\|e_{n}\|_{S(k)}\hskip 86.72377pt\forall n\in\mathbb{N},k\geq k_{0}.$
@@
$\|Tx\|_{k}=\bigg{\|}\sum^{\infty}_{n=1}x_{n}e_{n}\bigg{\|}\leq\sum^{\infty}_{n=1}|x_{n}|\|Te_{n}\|_{k}\leq C\sum^{\infty}_{n=1}|x_{n}|\|e_{n}\|_{S(k)}\leq C\|x\|_{S(k)}.$
This says that $\mathcal{A}$ is tame. ∎
In this paper, we will call an operator which is defined between Köthe spaces as a Toeplitz operator if its matrix is a Toeplitz matrix defined with respect to the Schauder basis $(e_{n})_{n\in\mathbb{N}}$. In sections 3 and 4, we will concentrate on lower triangular Toeplitz matrices and upper triangular Toeplitz matrices, respectively. We will determine the continuity and compactness of the operators whose associated matrix is related to these matrices. By collecting the results obtained in sections 3 and 4, we share the results about the continuity and compactness of a Toeplitz operator defined between power series spaces.
@@ 3. Operators Defined by A Lower Triangular Toeplitz Matrix
$\begin{pmatrix}\theta_{0}&0&0&0&\cdots\\ \theta_{1}&\theta_{0}&0&0&\cdots\\ \theta_{2}&\theta_{1}&\theta_{0}&0&\cdots\\ \theta_{3}&\theta_{2}&\theta_{1}&\theta_{0}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix}$
[…]
@@
In this section, we examine the properties of the operators defined by a lower triangular Toeplitz matrix between Köthe spaces. In [8], you can find the same idea for Proposition 3.1 and a generalized form Theorem 3.2 for $G_{1}$-spaces, arguing as in Theorem 2.1.
@@
Let $K(a_{n,k})$ and $K(b_{n,k})$ be Köthe spaces. If the operator $\widehat{T}_{\theta}:K(a_{n,k})\to K(b_{n,k})$ is continuous, then $\theta\in K(b_{n,k})$ and the following holds
$\quad\quad\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad\quad\quad b_{n,k}\leq Ca_{n,m}\quad\quad\quad\quad\forall n\in\mathbb{N}.$
In this case we can say that $K(a_{n,k})\subseteq K(b_{n,k})$.
@@
$\widehat{T}_{\theta}e_{n}=(0,\cdots,0,\theta_{0},\theta_{1},\theta_{2},\cdots)=\sum^{\infty}_{j=n}\theta_{j-n}e_{j}.$
@@
$\quad\quad\quad\quad\quad\|\widehat{T}_{\theta}e_{n}\|_{k}=\sum^{\infty}_{j=n}|\theta_{j-n}|b_{j,k}\leq C_{1}\|e_{n}\|_{m}=C_{1}a_{n,m}\quad\quad\quad\quad\forall n\in\mathbb{N}.$
Then, for all n and $j\geq n$, we have $|\theta_{j-n}|b_{j,k}\leq C_{1}a_{n,m}.$ By assuming $\theta_{0}\neq 0$, we can write that $b_{n,k}\leq C_{2}a_{n,m}$ for some $C_{2}>0$. This completes the proof. ∎
@@
Let $K(a_{n,k})$ be a Köthe space and $\Lambda_{1}(\beta)$ be a power series space of finite type. $\widehat{T}_{\theta}:K(a_{n,k})\to\Lambda_{1}(\beta)$ is continuous if and only if $\theta\in\Lambda_{1}(\beta)$ and the following condition holds:
$\quad\quad\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad\quad e^{-\frac{1}{k}\beta_{n}}\leq Ca_{n,m}\quad\quad\forall n\in\mathbb{N}.$ (3.1)
@@
The sufficiency part comes from Proposition 3.1. For the necessary part, assume $\theta\in\Lambda_{1}(\beta)$. By using 3.1 we have the following inequalities: for every $k\in\mathbb{N}$, there exist $m\in\mathbb{N}$ and $C_{1},C_{2}>0$ such that
$\begin{split}\|\widehat{T}_{\theta}e_{n}\|_{k}\leq\|\widehat{T}_{\theta}e_{n}\|_{2k}&=\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{2k}\beta_{j}}=\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{k}\beta_{j}}e^{-\frac{1}{k}\beta_{j}}\\ &\leq\bigg{(}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{k}\beta_{j}}\bigg{)}e^{-\frac{1}{k}\beta_{n}}\\ &\leq C_{1}\bigg{(}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{k}\beta_{j}}\bigg{)}a_{n,m}\\ &\leq C_{1}\bigg{(}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{k}\beta_{j-n+1}}\bigg{)}\|e_{n}\|_{m}\\ &\leq C_{1}\|\theta\|_{k}\|e_{n}\|_{m}\leq C_{2}\|e_{n}\|_{m}\end{split}$
Therefore, for all $k\in\mathbb{N}$, there exist $m\in\mathbb{N}$ such that
@@
that is, $\widehat{T}_{\theta}:K(a_{n,k})\to\Lambda_{1}(\beta)$ is continuous. ∎
@@
Let $K(a_{n,k})$ be a Köthe space and $\Lambda_{1}(\beta)$ be a nuclear power series space of finite type. $\widehat{T}_{\theta}:K(a_{n,k})\to\Lambda_{1}(\beta)$ is compact if and only if $\theta\in\Lambda_{1}(\beta)$ and the following condition holds:
$\quad\quad\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad e^{-\frac{1}{k}\beta_{n}}\leq Ca_{n,m}\quad\quad\forall n\in\mathbb{N}.$ (3.2)

### v3

@@ Abstract.
The aim of this paper is to define Toeplitz operators between Köthe spaces, especially power series spaces. We determine the conditions for continuity and compactness of these operators. We define the concept of S-tameness of a family of continuous operators. We construct some conditions on $S$-tameness for the families consisting of Toeplitz operators.
@@ 1. Introduction
The theory of Toeplitz operators defined on a Hilbert space such as the Hardy and the Bergman spaces is well-studied. The significance of this theory lies in the connection between operator theory and function spaces. The matrix of a Toeplitz operator defined on Hardy space of unit disk $H^{2}(\mathbb{D})$ is a Toeplitz matrix. Moreover, any bounded operator on $H^{2}(\mathbb{D})$ whose associated matrix is a Toeplitz matrix is a Toeplitz operator. We direct the reader to [8] for more information about Toeplitz operators defined on $H^{2}(\mathbb{D})$. In recent years, Toeplitz operators, whose ”associated” matrix is Toeplitz, are defined for more general topological vector spaces. For instance, in [3], Domański and Jasiczak developed the analogous theory for the space of $\mathcal{A}(\mathbb{R})$ real analytic functions on the real line. This space is not a Banach space, even not a metrizable space. In [6], Jasiczak introduced and characterized the class of Toeplitz operators on the Fréchet space of all entire functions $\mathcal{O}(\mathbb{C})$.
In [6], Jasiczak defined a continuous linear operator on $\mathcal{O}(\mathbb{C})$ as a Toeplitz operator if its matrix is a Toeplitz matrix. The matrix of an operator is defined with respect to the Schauder basis $(z^{n})_{n\in{N_{0}}}$. In this case, the symbol of a Toeplitz operator comes from the space $\mathcal{O}(\mathbb{C})\oplus(\mathcal{O}(\mathbb{C}))^{\prime}_{b}$ here $(\mathcal{O}(\mathbb{C}))^{\prime}_{b}$ is the strong dual of $\mathcal{O}(\mathbb{C})$. The space of entire functions $\mathcal{O}(\mathbb{C})$ is isomorphic to a power series space of infinite type $\Lambda_{\infty}(n)$. By taking inspiration from Jasiczak paper [6], we will define Toeplitz operators on more general power series spaces of finite or infinite type. Similarly, we will show that the symbol space is $\Lambda_{r}(\alpha)\oplus(\Lambda_{r}(\alpha))^{\prime}$, for $r=1,\infty$. We will also search the compactness of these operators and mention the conditions for the tameness of the families consisting of Toeplitz operators.
@@ 2. Preliminaries
In this section, we provide some fundamental facts and definitions essential for the subsequent discussions. We will use the standard terminology and notation of [9].
A complete Hausdorff locally convex space E whose topology defined by countable fundamental system of seminorms $(\|\cdot\|_{k})_{k\in\mathbb{N}}$ is called a Fréchet space. A matrix $(a_{n,k})_{k,n\in\mathbb{N}}$ of non-negative numbers is called a Köthe matrix if it is satisfies the following conditions:
@@
is called a Köthe space. Every Köthe space is a Fréchet space. From Proposition 27.13 of [9], the dual space of a Köthe space is
$(K(a_{n,k}))^{\prime}=\bigg{\{}y=(y_{n})_{n\in\mathbb{N}}\;\bigg{|}\;\sup_{n\in\mathbb{N}}|y_{n}a_{n,k}^{-1}|<+\infty\;\text{for some}\;k\in\mathbb{N}\bigg{\}}.$
By Grothendieck-Pietsch Criteria [9][Theorem 28.15] a Köthe space $K(a_{n,k})$ is nuclear if and only if for every $k\in\mathbb{N}$, there exists a $l>k$ so that
@@
For a nuclear Köthe spaces, $\|x\|_{k}=\sup_{n\in\mathbb{N}}|x_{n}|a_{n,k}$, $k\in\mathbb{N}$ forms an equivalent system of seminorms to the fundamental system of seminorms $\|x\|_{k}=\sum^{\infty}_{n=1}|x_{n}|a_{n,k}$, $k\in\mathbb{N}$, see Proposition 28.16 of [9].
Dynin-Mitiagin Theorem [9][Theorem 28.12] states that if a nuclear Fréchet space E with the sequence of seminorms $(\|f_{n}\|)_{n\in\mathbb{N}}$ has a Schauder basis $(f_{n})_{n\in\mathbb{N}}$, then it is canonically isomorphic to a nuclear Köthe space defined by the matrix $(\|f_{n}\|_{k})_{n,k\in\mathbb{N}}$. So, nuclear Köthe spaces hold a significant place in theory of nuclear Fréchet space.
Let $\alpha=\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ be a non-negative increasing sequence with $\lim_{n\rightarrow\infty}\alpha_{n}=\infty$. A power series space of finite type is defined by
@@
[…]
where $\mathbb{D}$ is the unit disk in $\mathbb{C}$ and $d\in\mathbb{N}$.
Let $E$ and $F$ be Fréchet spaces. A linear map $T:E\to F$ is called continuous if for every $k\in\mathbb{N}$ there exists $p\in\mathbb{N}$ and $C_{k}>0$ such that
@@
for all $x\in E$. A linear map $T:E\to F$ is called compact if $T(U)$ is precompact in $F$ where $U$ is a neighborhood of zero of E.
In this paper, we fixed the symbol $e_{n}$ to denote the sequence
@@
where 1 is in the nth place and 0 is in the others.
We will use the following Lemma to determine the continuity and compactness of operators defined between Köthe spaces.
@@
    If $K(b_{n,k})$ is Montel, then $T:K(a_{n,k})\to K(b_{n,k})$ is a compact operator if and only if there exists m such that for all k
@@
We want to note that a Fréchet space $E$ is Montel if each bounded set in $E$ is relatively compact and every power series space is Montel, see Theorem 27.9 of [9].
In the next proposition, it will be shown that the continuity condition is sufficient to ensure that linear operators defined only on the basis elements are well-defined.
Proposition 2.2.
Let $K(a_{n,k})$, $K(b_{n,k})$ be Köthe spaces and $(a_{n})_{n\in\mathbb{N}}\in K(b_{n,k})$ be a sequence. Let us define a linear map $T:K(a_{n,k})\to K(b_{n,k})$ such as
$Te_{n}=a_{n}\hskip 36.135pt\text{and}\hskip 36.135ptTx=\sum^{\infty}_{n=1}x_{n}Te_{n}$
for every $n\in\mathbb{N}$ and $x=\sum^{\infty}_{n=1}x_{n}e_{n}$. If the continuity condition
$\forall k\in\mathbb{N}\quad\quad\exists m\in\mathbb{N}\quad\quad\quad\sup_{n\in\mathbb{N}}\frac{\|Te_{n}\|_{k}}{\|e_{n}\|_{m}}<\infty$
holds, then $T$ is well-defined and continuous operator.
Proof.
Let $x=\sum^{\infty}_{n=1}x_{n}e_{n}$ be an arbitrary element of $K(a_{n,k})$ and assume that for every $k\in\mathbb{N}$ there exists a $m\in\mathbb{N}$ such that
$\sup_{n\in\mathbb{N}}\frac{\|Te_{n}\|_{k}}{\|e_{n}\|_{m}}<\infty.$
Now we define the sequence
$y_{N}=\sum^{N}_{n=1}T(x_{n}e_{n})=\sum_{n=1}^{N}x_{n}Te_{n}$
for every $N\in\mathbb{N}$. Then for every $k\in\mathbb{N}$, there exist $m\in\mathbb{N}$ and $C>0$ such that for sufficiently large $N,M\in\mathbb{N}$, $N>M$,
$\|y_{N}-y_{M}\|_{k}=\|\sum^{N}_{l=M+1}x_{l}T(e_{l})\|_{k}\leq\sum^{N}_{l=M+1}|x_{l}|\|e_{l}\|_{m}\frac{\|Te_{l}\|_{k}}{\|e_{l}\|_{m}}\leq C\sum^{N}_{l=M+1}|x_{l}|\|e_{l}\|_{m}$
which is arbitrarily small, since $\|x\|_{m}=\sum^{\infty}_{n=1}|x_{n}|\|e_{n}\|_{m}<\infty.$ This says that the sequence $(y_{N})_{N\in\mathbb{N}}\subseteq K(b_{k,n})$ is a Cauchy sequence and therefore has a limit. Let us say the $\lim_{N\to\infty}y_{N}=y$. Hence, we have
$y=\lim_{N\to\infty}y_{N}=\sum^{\infty}_{n=1}x_{n}T(e_{n})=T(x)\in K(b_{k,n}).$
Therefore $T$ is well-defined from $K(a_{k,n})$ to $K(b_{k,n})$ and $T$ is continuous from Lemma 2.1. ∎
@@
for each $x\in E$ and which define the topology. Every Fréchet space admits a grading. A graded Fréchet space is Fréchet space with a choice of grading. For more information, see [5]. In this paper, unless stated otherwise, we will assume that all Fréchet spaces are graded Fréchet spaces.
A pair of graded Fréchet spaces $(E,F)$ is called tame provided that there exists an increasing function $\sigma:\mathbb{N}\to\mathbb{N}$, such that for every continuous linear operator T from $E$ to $F$, there exists an $N$ and $C>0$ such that
$\hskip 83.11005pt\|Tx\|_{n}\leq C\|x\|_{\sigma(n)}\hskip 75.88371pt\forall x\in E\;\;\text{and}\;\;n\geq N.$
A Fréchet space $E$ is called tame if $(E,E)$ is tame. Tameness gives us a kind of control on the continuity of the operators. In [4], Dubinsky and Vogt used this notion to find a basis in complemented subspaces of some infinite type power series spaces. We refer to [1] and [7] for the other studies about tameness.
In this paper, instead of all operators defined on a Fréchet space, we will restrict ourselves to a subfamily of operators. With this aim, we will give the following definition.
@@
$\hskip 119.24506pt\|Tx\|_{k}\leq C\|x\|_{S(k)}\hskip 72.26999pt\forall x\in E,k\geq k_{0}.$
@@
Let $K(a_{n,k})$ and $K(b_{n,k})$ be Köthe spaces and $S:\mathbb{N}\to\mathbb{N}$ be a non-decreasing function. A family of linear continuous operators $\mathcal{A}\subseteq L(K(a_{n,k}),K(b_{n,k}))$ is S-tame if and only if for every operator $T\in\mathcal{A}$ there exist $k_{0}\in\mathbb{N}$ and $C>0$ such that
$\hskip 101.17755pt\|Te_{n}\|_{k}\leq C\|e_{n}\|_{S(k)}\hskip 79.49744pt\forall n\in\mathbb{N},k\geq k_{0}.$
@@
Let us assume that for every $T\in\mathcal{A}$ there exist a $k_{0}\in\mathbb{N}$ and $C>0$ such that
$\hskip 101.17755pt\|Te_{n}\|_{k}\leq C\|e_{n}\|_{S(k)}\hskip 79.49744pt\forall n\in\mathbb{N},k\geq k_{0}.$
@@
$\begin{split}\|Tx\|_{k}&=\bigg{\|}\sum^{\infty}_{n=1}x_{n}T(e_{n})\bigg{\|}_{k}\leq\sum^{\infty}_{n=1}|x_{n}|\|Te_{n}\|_{k}\leq C\sum^{\infty}_{n=1}|x_{n}|\|e_{n}\|_{S(k)}\\ &=C\sum^{\infty}_{n=1}|x_{n}|a_{n,S(k)}=C\|x\|_{S(k)}.\end{split}$
This says that $\mathcal{A}$ is S-tame. The other direction is straightforward. ∎
In this paper, we will call an operator which is defined between Köthe spaces as a Toeplitz operator if its matrix is a Toeplitz matrix defined with respect to the Schauder basis $(e_{n})_{n\in\mathbb{N}}$. In sections 3 and 4, we will concentrate on lower triangular Toeplitz matrices and upper triangular Toeplitz matrices, respectively. We will determine the continuity and compactness of the operators whose associated matrix is related to these matrices. By collecting the results obtained in sections 3 and 4, we share the results about the continuity and compactness of a Toeplitz operator defined between power series spaces in section 5.
@@ 3. Operators Defined by A Lower Triangular Toeplitz Matrix
$\begin{pmatrix}\theta_{0}&0&0&0&\cdots\\ \theta_{1}&\theta_{0}&0&0&\cdots\\ \theta_{2}&\theta_{1}&\theta_{0}&0&\cdots\\ \theta_{3}&\theta_{2}&\theta_{1}&\theta_{0}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix}.$
We want to define an operator $\widehat{T}_{\theta}:K(a_{n,k})\to K(b_{n,k})$ by taking $\widehat{T}_{\theta}e_{n}$ as the nth column of the above matrix, that is,
@@
In fact, we can not guarantee that the operator $\widehat{T}_{\theta}$ is correctly defined between two general Köthe spaces $K(a_{n,k})$ and $K(b_{n,k})$, because we do not know if the series $\sum^{\infty}_{n=1}x_{n}\widehat{T}_{\theta}e_{n}$ converges in $K(b_{n,k})$ for every $x\in K(a_{n,k})$ in this general case. Below, we will examine the cases where this operator can be correctly defined, and in these cases, we will analyze the continuity and compactness of this operator. Throughout this paper, we will assume that the sequences $\theta$ satisfy the following condition
$\forall n\in\mathbb{N}\hskip 14.45377pt\exists s>n\hskip 43.36243pt\theta_{s}\neq 0.$
Otherwise, $\theta$ produces a finite rank operator, which is obviously continuous and compact.
The idea of Proposition 3.1 is the same as [10, Theorem 2.1]. Therein one can also find a generalized form of Theorem 3.3.
@@
Let $K(a_{n,k})$, $K(b_{n,k})$ be Köthe spaces and assume that $\theta$ be a sequence and $s_{0}=\min\{t:\theta_{t}\neq 0\}$. If $\widehat{T}_{\theta}:K(a_{n,k})\to K(b_{n,k})$ is a continuous operator, then $\theta\in K(b_{n,k})$ and the following holds
$\hskip 43.36243pt\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad b_{n+s_{0},k}\leq Ca_{n,m}\hskip 36.135pt\forall n\in\mathbb{N}.$
@@
$\widehat{T}_{\theta}e_{n}=(0,\dots,0,\theta_{0},\theta_{1},\theta_{2},\dots)=\sum^{\infty}_{j=n}\theta_{j-n}e_{j}\in K(b_{n,k}).$
@@
$\hskip 50.58878pt\|\widehat{T}_{\theta}e_{n}\|_{k}=\sum^{\infty}_{j=n}|\theta_{j-n}|b_{j,k}\leq C_{1}\|e_{n}\|_{m}=C_{1}a_{n,m}\hskip 50.58878pt\forall n\in\mathbb{N}.$
Then, for all n and $j\geq n$, we have $|\theta_{j-n}|b_{j,k}\leq C_{1}a_{n,m}.$ Hence we can write that $b_{n+s_{0},k}\leq C_{2}a_{n,m}$ for some $C_{2}>0$. This completes the proof. ∎
For compact $\widehat{T}_{\theta}:K(a_{n,k})\to K(b_{n,k})$ operators, the relationship between Köthe matrices $(a_{n,k})_{n,k\in\mathbb{N}}$ and $(b_{n,k})_{n,k\in\mathbb{N}}$ is as follows:
Proposition 3.2.
Let $K(a_{n,k})$ and $K(b_{n,k})$ be Köthe spaces such that $K(b_{n,k})$ is Montel and assume that $\theta\in K(b_{n,k})$ and $s_{0}=\min\{t:\theta_{t}\neq 0\}$. If $\widehat{T}_{\theta}:K(a_{n,k})\to K(b_{n,k})$ is a compact operator, then the following holds
$\hskip 36.135pt\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad b_{n+s_{0},k}\leq Ca_{n,m}\hskip 36.135pt\forall n\in\mathbb{N}.$
Proof.
The proof is similar to the proof of Proposition 3.1. ∎
@@
Let $K(a_{n,k})$ be a Köthe space, $\Lambda_{1}(\beta)$ be a power series space of finite type and assume that $\theta$ be a sequence and $s_{0}=\min\{t:\theta_{t}\neq 0\}$. $\widehat{T}_{\theta}:K(a_{n,k})\to\Lambda_{1}(\beta)$ is well-defined and continuous if and only if $\theta\in\Lambda_{1}(\beta)$ and the following condition holds:
$\hskip 14.45377pt\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad e^{-\frac{1}{k}\beta_{n+s_{0}}}\leq Ca_{n,m}\hskip 36.135pt\forall n\in\mathbb{N}.$ (3.1)
@@
If the operator $\widehat{T}_{\theta}$ is continuous, Proposition 3.1 gives us that $\theta\in\Lambda_{1}(\beta)$ and (3.1) holds. For the other direction, let us assume $\theta\in\Lambda_{1}(\beta)$. By using the condition (3.1) we have the following: for every $k\in\mathbb{N}$, there exist $m\in\mathbb{N}$ and $C_{1},C_{2}>0$ such that
$\begin{split}\|\widehat{T}_{\theta}e_{n}\|_{k}&\leq\|\widehat{T}_{\theta}e_{n}\|_{2k}=\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{2k}\beta_{j}}\\ &=\sum^{\infty}_{j=n+s_{0}}|\theta_{j-n}|e^{-\frac{1}{2k}\beta_{j}}e^{\frac{1}{k^{2}+2k}\beta_{n+s_{0}}}e^{-\frac{1}{k^{2}+2k}\beta_{n+s_{0}}}\\ &\leq C_{1}a_{n,m}\sum^{\infty}_{j=n+s_{0}}|\theta_{j-n}|e^{-\frac{1}{2k}\beta_{j}}e^{\frac{1}{k^{2}+2k}\beta_{n+s_{0}}}\\ &=C_{1}a_{n,m}\sum^{\infty}_{j=n+s_{0}}|\theta_{j-n}|e^{-\frac{1}{2(k+2)}\beta_{j}}e^{-\frac{1}{k^{2}+2k}(\beta_{j}-\beta_{n+s_{0}})}\\ &\leq C_{1}a_{n,m}\sum^{\infty}_{j=n+s_{0}}|\theta_{j-n}|e^{-\frac{1}{2(k+2)}\beta_{j}}=C_{1}\|\theta\|_{2(k+2)}a_{n,m}=C_{2}a_{n,m}\end{split}$
for every $n\in\mathbb{N}$. In the second line of the inequalites we use the fact that
$\frac{1}{k^{2}+2k}=\frac{1}{2k}-\frac{1}{2(k+2)}$
for every $k\in\mathbb{N}$. Therefore, $\widehat{T}_{\theta}e_{n}\in\Lambda_{1}(\beta)$ for every $n\in\mathbb{N}$ and for every $k\in\mathbb{N}$ there exist $m\in\mathbb{N}$ such that
@@
that is, $\widehat{T}_{\theta}:K(a_{n,k})\to\Lambda_{1}(\beta)$ is well defined and continuous by Proposition 2.2. ∎
@@
Let $K(a_{n,k})$ be a Köthe space, $\Lambda_{1}(\beta)$ be a power series space of finite type and assume that $\theta$ be a sequence and $s_{0}=\min\{t:\theta_{t}\neq 0\}$. $\widehat{T}_{\theta}:K(a_{n,k})\to\Lambda_{1}(\beta)$ is compact if and only if $\theta\in\Lambda_{1}(\beta)$ and the following condition holds:
[…]